OpenStudy (anonymous):

evaluate the integral ∫xsin x/2 dx

OpenStudy (anonymous):

parts?

OpenStudy (anonymous):

yeah integrate by parts. \[\int fdg=fg-\int gdf\] use \[f=x\] \[df=1\] \[dg=\sin(\frac{x}{2})\] \[g=-2\cos(\frac{x}{2})\] and you get it in one step basically \[-2x\cos(x)+2\int\cos(\frac{x}{2})dx\]